a particle sliding down an inclined plane travels 2 feet the first second. In any second after the first it slides 4 feet farther than it did in the preceding second. how far will it slide in 7th second? how far will it slide in 7 seconds? how long will it take to slide 128 feet?
Answer by mananth(16946)

a particle sliding down an inclined plane travels 2 feet the first second. In any second after the first it slides 4 feet farther than it did in the preceding second. how far will it slide in 7th second? how far will it slide in 7 seconds? how long will it take to slide 128 feet?

t1=2 =a
t2=2+4=6
t3=6+4=10
a=2
d=4
tn= a+(n-1)d
t7=2+(7-1)4
t7= 26, 26 feet in the 7th. second
Sn=n/2[2a+(n-1)d]
S7=7/2[2*2+(7-1)4]
S7=7/2[28]
S7=98, 98 feet in 7 seconds

128=n/2[4+(n-1)2]
256=4n
256=4n +2n^2 -2n
2n^2+2n-256=0
/2
n^2+n-128=0
solve the quadratic for x
you will get n=10.82
In 10.82 sec it will go 128 feet